OpenStudy (anonymous):

Solve for x: \[7+2lnx=6\]

OpenStudy (australopithecus):

2ln(x) = 6 - 7

OpenStudy (anonymous):

-2 ln(x)= -6 ln(x)= 3 e^(ln(x))= e^3 x=e^3, Hope this helps (✿◠‿◠) .

OpenStudy (australopithecus):

ln(x) = -1/2 e^(ln(x)) = e^(-1/2) x = e^(-1/2) x = 1/e^(1/2)

OpenStudy (australopithecus):

i think you made a mistake in your addition

OpenStudy (australopithecus):

@TheOlympicnomad

OpenStudy (australopithecus):

Yes you are wrong https://www.wolframalpha.com/input/?i=7%2B2lnx%3D6

OpenStudy (anonymous):

7 + 2lnx = 6 2lnx = -1 lnx = -1/2 x = e^(-1/2)

OpenStudy (australopithecus):

remember \[\ln(x) = \log_e(x)\]

OpenStudy (australopithecus):

when you make a logarithm a power to its base the logarithm cancels out and you are left with what was inside it

OpenStudy (australopithecus):

we just write it as ln because it is so commonly used people are lazy

OpenStudy (anonymous):

(►.◄)!!.